OpenStudy (barrelracing):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3. A) -91 b) -33 c) -25 d) 35

OpenStudy (michele_laino):

hint, by definition of compiste function, we can write this: \[\Large f\left( {g\left( x \right)} \right) = 2g\left( x \right) + 1 = 2\left( { - 5x + 2} \right) + 1 = ...?\]

OpenStudy (michele_laino):

please continue

OpenStudy (barrelracing):

do I have to plug something in or just solve?

OpenStudy (michele_laino):

solve as an algebraic expression

OpenStudy (michele_laino):

hint: \[\Large 2\left( { - 5x + 2} \right) + 1 = - 10x + 4 + 1 = ...?\]

OpenStudy (barrelracing):

it all cancels out. -10x + 4 + 1 = -10x + 4 + 1

OpenStudy (michele_laino):

now simplify further, what do you get?

OpenStudy (michele_laino):

4+1=...?

OpenStudy (barrelracing):

5!!! :)

OpenStudy (michele_laino):

ok! so we got this: \[\Large f\left( {g\left( x \right)} \right) = - 10x + 5\]

OpenStudy (michele_laino):

please substitute x=3 at the right side, what do you get?

OpenStudy (barrelracing):

f(g(x)) = -25

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

\[f\left( {g\left( 3 \right)} \right) = - 25\]

OpenStudy (barrelracing):

thank you so much for your help

OpenStudy (michele_laino):

:)